A bird watcher meanders through the woods, walking 0.45 km due east, 0.60 km due south, and 2.85 km in a direction 41.0° north of west. The time required for this trip is 2.50 h.

(a) Determine the magnitude and direction (relative to due west) of the bird watcher's displacement. Use kilometers and hours for distance and time, respectively.

magnitude km
direction ° north of west


(b) Determine the magnitude and direction (relative to due west) of the bird watcher's average velocity.

magnitude km/hr
direction ° north of west
Answer #1

a)

X-component of displacement

X=0.45-2.85Cos(41)=-1.7 km

Y-Component of displacement

Y=-0.6+2.85Sin(41)=1.27 km

Magnitude

r=sqrt[(-1.7)2+1.272] =2.122 km

Direction

o=tan-1(1.27/01.7) =-36.76o =36.76o North of west

b)

Vavg=r/t =2.122/2.5

Vavg=0.8487 Km/hr

Direction

o=36.76o North of west